Find the whole quantity if 
70% of it is 14 minutes
Find the whole quantity if 
70% of it is 14 minutes

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Let us assume the whole quantity be x, Then,
(70/100) × (x) = 14
X = 14 × (100/70)
X = 14 × (10/7)
X = 20 minutes

Thanks